# NRGBoost: Energy-Based Generative Boosted Trees

### v1

@@ Abstract
Despite the rise to dominance of deep learning in unstructured data domains, tree-based methods such as Random Forests (RF) and Gradient Boosted Decision Trees (GBDT) are still the workhorses for handling discriminative tasks on tabular data. We explore generative extensions of these popular algorithms with a focus on explicitly modeling the data density (up to a normalization constant), thus enabling other applications besides sampling. As our main contribution we propose an energy-based generative boosting algorithm that is analogous to the second order boosting implemented in popular packages like XGBoost. We show that, despite producing a generative model capable of handling inference tasks over any input variable, our proposed algorithm can achieve similar discriminative performance to GBDT on a number of real world tabular datasets, outperforming alternative generative approaches. At the same time, we show that it is also competitive with neural network based models for sampling.
@@ 1 Introduction
Generative models have achieved tremendous success in computer vision and natural language processing, where the ability to generate synthetic data guided by user prompts opens up many exciting possibilities. While generating synthetic table records does not necessarily enjoy the same wide appeal, this problem has still received considerable attention as a potential avenue for bypassing privacy concerns when sharing data. Estimating the data density, $p(\mathbf{x})$, is another typical application of generative models which enables a host of different use cases that can be particularly interesting for tabular data. Unlike discriminative models which are trained to perform inference over a single target variable, density models can be used more flexibly for inference over different variables or for out of distribution detection. They can also handle inference with missing data in a principled way by marginalizing over unobserved variables.
@@
The development of generative models for tabular data has mirrored its progression in computer vision with many of its Deep Learning (DL) approaches being adapted to the tabular domain (Jordon et al., 2018; Xu et al., 2019; Fan et al., 2020; Engelmann & Lessmann, 2021; Zhao et al., 2021; Kotelnikov et al., 2023). Unfortunately, these methods are only useful for sampling as they either don’t model the density explicitly or can’t evaluate it due to untractable marginalization over high dimensional latent variable spaces. Furthermore, despite growing in popularity, DL has still failed to displace tree-based ensemble methods as the tool of choice for handling tabular discriminative tasks with gradient boosting still being found to outperform neural-network-based methods in many real world datasets (Grinsztajn et al., 2022; Borisov et al., 2022a).
@@
While there have been recent efforts to extend the success of tree-based models to generative modeling (Correia et al., 2020; Wen & Hang, 2022; Nock & Guillame-Bert, 2022; Watson et al., 2023; Nock & Guillame-Bert, 2023; Jolicoeur-Martineau et al., 2024), direct extensions of Random Forests (RF) and Gradient Boosted Decision Trees (GBDT) are still missing. It is this gap that we try to address, seeking to keep the general algorithmic structure of these popular algorithms but replacing the optimization of their discriminative objective with a generative counterpart. Our main contributions in this regard are:
@@
  • We propose NRGBoost, a novel energy-based generative boosting model that is trained to maximize a local second order approximation to the likelihood at each boosting round.
@@
  • We propose an amortized sampling approach that greatly reduces the training time of NRGBoost, which, like other energy-based models, is often dominated by sampling.
@@
  • We explore the use of bagged ensembles of Density Estimation Trees (DET) (Ram & Gray, 2011) with feature subsampling as a generative counterpart to Random Forests.
@@
The longstanding popularity of GBDT models in machine learning practice can, in part, be attributed to the strength of its empirical results and the efficiency of its existing implementations. We therefore focus on an experimental evaluation in real world datasets spanning a range of use cases, number of samples and features. On smaller datasets, our implementation of NRGBoost can be trained in a few minutes on a mid-range consumer CPU and achieves similar discriminative performance to a standard GBDT model, while remaining competitive with the best generative models for sampling.
Figure 1: Downsampled MNIST samples generated by the best generative models on this dataset. Despite this being a simple image dataset, it can be challenging for tabular generative models due to the high dimensionality and complex structure of correlations between features. We find NRGBoost to be the only model that is able to generate passable samples.
@@ 2 Energy Based Models
Here $f(\mathbf{x}):\mathcal{X}\rightarrow\mathbb{R}$ is a real function over the input domain.111 We will assume that $\mathcal{X}$ is finite and discrete to simplify the notation and exposition but everything is applicable to bounded continuous input spaces, replacing the sums with integrals as appropriate. We will avoid introducing any parametrization, instead treating the function $f\in\mathcal{F}(\mathcal{X})$ lying in an appropriate function space over the input space as our model parameter directly. $Z[f]=\sum_{\mathbf{x}\in\mathcal{X}}\exp\left(f(\mathbf{x})\right)$, known as the partition function, is then a functional of $f$ giving us the necessary normalizing constant.
This is the most flexible way one could represent a probability density function making essentially no compromises on its structure. The downside to this is that for most interesting choices of $\mathcal{F}$, computing or estimating this normalizing constant is untractable which makes training these models difficult. Their unnormalized nature, however, does not prevent EBMs from being useful in a number of applications besides sampling. They can be used to perform inference over small enough groups of variables. Partitioning the input space into a set of observed variables $\mathbf{x}_{o}$ and unobserved variables $\mathbf{x}_{u}$ we have
@@
which only involves normalizing (i.e., computing a softmax) over the possible values of the unobserved variables. Furthermore, for anomaly or out of distribution detection, knowledge of the normalizing constant is not necessary.
One common way to train an energy-based model to approximate a data generating distribution, $p(\mathbf{x})$, is to minimize the Kullback-Leibler divergence between $p$ and $q_{f}$, or equivalently, maximize the expected log likelihood functional:
@@
This optimization is typically carried out by gradient descent over the parameters of $f$, but due to the intractability of the partition function, one must rely on Markov Chain Monte Carlo (MCMC) sampling to estimate the gradients (Song & Kingma, 2021).
@@ 3 NRGBoost
Expanding the increase in log-likelihood in equation 3 due to a variation $\delta f$ around an energy function $f$ up to second order we have
@@
The $\delta f$ that maximizes this quadratic approximation should thus have a large positive difference between the expected value under the data and under $q_{f}$ while having low variance under $q_{f}$. We note that just like the original log-likelihood, this Taylor expansion is invariant to adding an overall constant to $\delta f$. This means that, in maximizing equation 4 we can consider only functions that have zero expectation under $q_{f}$ in which case we can simplify $\Delta L_{f}[\delta f]$ as
@@
We thus formulate our boosting algorithm as modelling the data density with an additive energy function. At each boosting iteration we improve upon the current energy function $f_{t}$ by finding an optimal step $\delta f^{*}_{t}$ that maximizes $\Delta L_{f_{t}}[\delta f]$:
$\delta f^{*}_{t}=\arg\max_{\delta f\in\mathcal{H}_{t}}\Delta L_{f_{t}}[\delta f]\,,$ (6)
where $\mathcal{H}_{t}$ is an appropriate space of functions (satisfying $\mathbb{E}_{\mathbf{x}\sim q_{f_{t}}}\delta f(\mathbf{x})=0$ if equation 5 is used). The solution to this problem can be interpreted as a Newton step in the space of energy functions. Because for an EBM, the Fisher Information matrix with respect to the energy function and the hessian of the expected log-likelihood are the same (see Appendix A), we can also interpret the solution to equation 6 as a natural gradient step. This approach is analogous to the second order step implemented in modern gradient boosting libraries such as XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017) and which can be traced back to Friedman et al. (2000).
In updating the current iterate, $f_{t+1}=f_{t}+\alpha_{t}\cdot\delta f^{*}_{t}$, we scale $\delta f^{*}_{t}$ by an additional scalar step-size $\alpha_{t}$. This can be interpreted as a globalization strategy to account for the fact that the quadratic approximation in equation 4 is not necessarily valid over large steps in function space. A common strategy in nonlinear optimization would be to select $\alpha_{t}$ via a line search based on the original log-likelihood. Common practice in discriminative boosting however is to interpret this step size as a regularization parameter and to select a fixed value in ]0,1]]0,1]] 0 , 1 ] with (more) smaller steps typically outperforming fewer larger ones when it comes to generalization. We choose to adopt a hybrid strategy, first selecting an optimal step size by line search and then shrinking it by a fixed factor. We find that this typically accelerates convergence allowing the algorithm to take comparatively larger steps that increase the likelihood in the initial phase of boosting.
For a starting point, $f_{0}$, we can choose the logarithm of any probability distribution over $\mathcal{X}$ as long as it is easy to evaluate. Sensible choices are a uniform distribution (i.e., $f\equiv 0$), the product of marginals for the training set, or any mixture distribution between these two. In Figure 2 we show an example of NRGBoost starting from a uniform distribution and learning a toy 2D data density.
@@ 3.1 Weak Learners
Figure 2: Density learned by NRGBoost at different boosting iterations (1, 3, 10 and 100), for a toy dataset (shown right), starting from a uniform distribution. Weak learners are trees with 16 leaves.
As a weak learner we will consider piecewise constant functions defined by binary trees over the input space. Letting $\bigcup_{j=1}^{J}X_{j}=\mathcal{X}$ be the partitioning of the input space induced by the leaves of a binary tree whose internal nodes represent a split along one dimension into two disjoint partitions, we take as $\mathcal{H}$ the set of functions such as
$\delta f(\mathbf{x})=\sum_{j=1}^{J}w_{j}\mathbf{1}_{X_{j}}(\mathbf{x})\,,$ (7)
where $\mathbf{1}_{X}$ denotes the indicator function of a subset $X$ and $w_{j}$ are values associated with each leaf $j\in[1..J]$. In a standard decision tree these values would typically encode an estimate of $p(y|\mathbf{x}\in X_{j})$, with $y$ being a special target variable that is never considered for splitting. In our generative approach they encode unconditional densities (or more precisely energies) over each leaf’s support and every variable can be used for splitting. Note that our functions $\delta f$ are thus parametrized by the values $w_{j}$ as well the structure of the tree and the variables and values for the split at each node which ultimately determine the $X_{j}$. We omit these dependencies for brevity.
Replacing the definition in equation 7 in our objective (equation 5) we get the following optimization problem to find the optimal decision tree:
@@
where $P(X_{j})$ and $Q_{f}(X_{j})$ denote the probability of the event $\mathbf{x}\in X_{j}$ under the respective distribution and the constraint ensures that $\delta f$ has zero expectation under $q_{f}$. With respect to the leaf weights this is a quadratic program whose optimal solution and objective values are respectively given by
@@
Because carrying out the maximization of this optimal value over the tree structure that determines the $X_{j}$ is hard, we approximate its solution by greedily growing a tree that maximizes it when considering how to split each node individually. A parent leaf with support $X_{P}$ is thus split into two child leaves, with disjoint support, $X_{L}\cup X_{R}=X_{P}$, so as to maximize over all possible partitionings along a single dimension, $\mathcal{P}\left(X_{P}\right)$, the following objective:
@@
Note that, when using parametric weak learners, computing a second order step would typically involve solving a linear system with a full Hessian. As we can see, this is not the case when the weak learners are decision trees, where the optimal value to assign to a leaf $j$ does not depend on any information from other leaves and, likewise, the optimal objective value is a sum of terms, each depending only on information from a single leaf. This would have not been the case had we tried to optimize the likelihood functional in Equation 3 directly instead of its quadratic approximation.
@@ 3.2 Amortized Sampling
To compute the leaf values in equation 9 and the splitting criterion in equation 10 we would have to know $P(X)$ and be able to compute $Q_{f}(X)$ which is infeasible due to the untractable normalization constant. We therefore estimate these quantities, with recourse to empirical data for $P(X)$, and to samples approximately drawn from the model with MCMC. Because even if the input space is not partially discrete, $f$ is still discontinuous and constant almost everywhere we can’t use gradient based samplers and therefore rely on Gibbs sampling instead. This only requires evaluating each $f_{t}$ along one dimension at a time, while keeping all others fixed which can be computed efficiently for a tree by traversing it only once. However, since at boosting iteration $t$ our energy function is a sum of $t$ trees, this computation scales linearly with the iteration number. This makes the overall time spent sampling quadratic in the number of boosting iterations and thus precluding us from training models with a large number of trees.
@@
In order to reduce the burden associated with this sampling, which dominates the runtime of training the model, we propose a new sampling approach that leverages the cumulative nature of boosting. The intuition behind this approach is that the set of samples used in the previous boosting round are (approximately) drawn from a distribution that is already close to the new model distribution. It could therefore be helpful to keep some of those samples, especially those that conform the best to the new model. Rejection sampling allows us to do just that. The boosting update in terms of the densities takes the following multiplicative form:
$q_{t}(\mathbf{x})=k_{t}\,q_{t-1}(\mathbf{x})\exp\left(\alpha_{t}\delta f_{t}(\mathbf{x})\right)\,.$ (11)
@@
Here, $k_{t}$ is an unknown multiplicative constant and since $\delta f_{t}$ is given by a tree, we can easily bound the exponential factor by finding the leaf with the largest value. We can therefore use the previous model, $q_{t-1}(\mathbf{x})$, as a proposal distribution for which we already have a set of samples and keep each sample, $\mathbf{x}$, with an acceptance probability of:
$p_{accept}(\mathbf{x})=\exp\left[\alpha_{t}\left(\delta f_{t}(\mathbf{x})-\max_{\mathbf{x}}\delta f_{t}(\mathbf{x})\right)\right]\,.$ (12)
We note that knowledge of the constant $k_{t}$ is not necessary to compute this probability.
@@
Our proposed sampling strategy is to maintain a fixed size pool of approximate samples from the model. At the end of each boosting round, we use rejection sampling to remove samples from the pool and draw new samples from the model using Gibbs sampling. Note that $q_{0}$ is typically a simple model for which we can both directly evaluate the desired quantities (i.e., $Q_{0}(X)$ for a given partition $X$) and cheaply draw exact samples from. As such, no samples are required for the first iteration of boosting and for the second we can initialize the sample pool by drawing exact samples from $q_{1}$ with rejection sampling using $q_{0}$ as a proposal distribution.
@@
This approach works better when the range of $\delta f_{t}$ and/or the step sizes $\alpha_{t}$ are small as this leads to larger acceptance probabilities. Note that in practice it can be helpful to independently refresh a fixed fraction of samples, $p_{refresh}$, at each round of boosting in order to encourage more diverse samples between rounds. This can be accomplished by keeping each sample with a probability $p_{accept}(\mathbf{x})(1-p_{refresh})$ instead.

### v2

@@ Abstract
Despite the rise to dominance of deep learning in unstructured data domains, tree-based methods such as Random Forests (RF) and Gradient Boosted Decision Trees (GBDT) are still the workhorses for handling discriminative tasks on tabular data. We explore generative extensions of these popular algorithms with a focus on explicitly modeling the data density (up to a normalization constant), thus enabling other applications besides sampling. As our main contribution we propose an energy-based generative boosting algorithm that is analogous to the second-order boosting implemented in popular libraries like XGBoost. We show that, despite producing a generative model capable of handling inference tasks over any input variable, our proposed algorithm can achieve similar discriminative performance to GBDT on a number of real world tabular datasets, outperforming alternative generative approaches. At the same time, we show that it is also competitive with neural-network-based models for sampling. Code is available at https://github.com/ajoo/nrgboost.
@@ 1 Introduction
Generative models have achieved tremendous success in computer vision and natural language processing, where the ability to generate synthetic data guided by user prompts opens up many exciting possibilities. While generating synthetic table records does not necessarily enjoy the same wide appeal, this problem has still received considerable attention as a potential avenue for bypassing privacy concerns when sharing data. Estimating the data density, $p(\mathbf{x})$, is another typical application of generative models which enables a host of different use cases that can be particularly interesting for tabular data. Unlike discriminative models, which are trained to perform inference over a single target variable, density models can be used more flexibly for inference over different variables or for out-of-distribution detection. They can also handle inference with missing data in a principled way by marginalizing over unobserved variables.
@@
The development of generative models for tabular data has mirrored its progression in computer vision, with many of its deep learning approaches being adapted to the tabular domain (Jordon et al., 2018; Xu et al., 2019; Engelmann & Lessmann, 2021; Fan et al., 2020; Zhao et al., 2021; Kotelnikov et al., 2023). Unfortunately, these methods are only useful for sampling as they either do not model the density explicitly or can not evaluate it efficiently due to intractable marginalization over high-dimensional latent variable spaces. Furthermore, despite growing in popularity, deep learning has still failed to displace tree-based ensemble methods as the tool of choice for handling tabular discriminative tasks, with gradient boosting still being found to outperform neural-network-based methods in many real world datasets (Grinsztajn et al., 2022; Borisov et al., 2022a).
@@
While there have been recent efforts to extend the success of tree-based models to generative modeling (Correia et al., 2020; Wen & Hang, 2022; Nock & Guillame-Bert, 2022; Watson et al., 2023; Nock & Guillame-Bert, 2023; Jolicoeur-Martineau et al., 2024; McCarter, 2024), direct extensions of Random Forests (RF) and Gradient Boosted Decision Trees (GBDT) are still missing. We try to address this gap, seeking to keep the general algorithmic structure of these popular algorithms but replacing the optimization of their discriminative objective with a generative counterpart. Our main contributions in this regard are:
@@
  • We propose NRGBoost, a novel energy-based generative boosting model that is trained to maximize a local second-order approximation to the log-likelihood at each boosting round.
@@
  • We propose an amortized sampling approach that significantly reduces the training time of NRGBoost, which, like other energy-based models, is often dominated by the Markov chain Monte Carlo sampling required for training.
@@
  • We explore the use of bagged ensembles of Density Estimation Trees (Ram & Gray, 2011) with feature subsampling as a generative counterpart to Random Forests.
@@
The longstanding popularity of GBDT models in machine learning practice can, in part, be attributed to the strength of its empirical results and the efficiency of its existing implementations. We therefore focus on an experimental evaluation in real world datasets spanning a range of use cases, number of samples and features. On smaller datasets, our implementation of NRGBoost can be trained in a few minutes on a typical consumer CPU and achieves similar discriminative performance to a standard GBDT model, while remaining competitive with other generative models for sampling.
Figure 1: Overview of an NRGBoost training iteration: 1) Draw new samples from an ensemble of trees representing a current energy function, $f$, and add them to the sample pool. 2) Fit a new weak learner (tree) representing a model update, $\delta f$. 3) Update the model. 4) Use rejection sampling to discard samples from the sample pool that conform poorly to the new model.
@@ 2 Energy Based Models
Here, $f(\mathbf{x}):\mathcal{X}\rightarrow\mathbb{R}$ is a real function over the input domain.111 We will assume that $\mathcal{X}$ is finite and discrete to simplify the notation and exposition but everything is applicable to bounded continuous input spaces, replacing the sums with integrals as appropriate. We will avoid introducing any parametrization and instead treat the function $f\in\mathcal{F}(\mathcal{X})$, lying in an appropriate function space over the input domain, as our model parameter directly. $Z[f]=\sum_{\mathbf{x}\in\mathcal{X}}\exp\left(f(\mathbf{x})\right)$, known as the partition function, is then a functional of $f$ giving us the necessary normalizing constant.
This is the most flexible way one could represent a probability density function, making essentially no compromises on its structure. It can seamlessly handle input domains comprising a mixture of continuous and categorical features, making it a natural choice for tabular data. The downside to this is that, for most interesting choices of $\mathcal{F}$, computing the normalizing constant is intractable which makes training these models difficult. Their unnormalized nature, however, does not prevent EBMs from being useful in a number of applications besides sampling, such as performing joint inference over small enough groups of variables. For example, by splitting the input domain into a set of observed variables, $\mathbf{x}_{o}$, and unobserved variables, $\mathbf{x}_{u}$, we have
@@
which only involves normalizing (i.e., computing a softmax) over all of the possible values of the unobserved variables. EBMs can also handle inference with missing data in a principled manner by marginalizing over unobserved variables. Furthermore, for anomaly or out-of-distribution detection, knowledge of the normalizing constant is not necessary.
One common way to train an energy-based model to approximate a data generating distribution, $p(\mathbf{x})$, is to minimize the Kullback-Leibler divergence between $p$ and $q_{f}$, or equivalently, maximize the expected log-likelihood functional:
@@
This optimization is typically carried out by gradient descent over the parameters of $f$. However, due to the intractability of the partition function, one must rely on Markov chain Monte Carlo (MCMC) sampling to estimate the gradients (Song & Kingma, 2021).
@@ 3 NRGBoost
Expanding the increase in log-likelihood in Equation 3 due to a variation $\delta f$ around an energy function $f$ up to second order (see Appendix A), we have
@@
The $\delta f$ that maximizes this quadratic approximation should thus have a large positive difference between the expected value under the data and under $q_{f}$, while having low variance under $q_{f}$. Note that, just like the original log-likelihood, this Taylor expansion is invariant to adding a constant to $\delta f$ (i.e., $\Delta L_{f}[\delta f+c]=\Delta L_{f}[\delta f]$ for a constant function $c$). This means that, in maximizing Equation 4, we can restrict our attention to functions that have zero expectation under $q_{f}$, in which case we can simplify $\Delta L_{f}$ as
@@
We can thus formulate a boosting algorithm that, at each boosting iteration, $t$, improves upon a current energy function, $f_{t}$, by finding an optimal step, $\delta f^{*}_{t}$, that maximizes $\Delta L_{f_{t}}$:
$\delta f^{*}_{t}=\operatorname*{arg\,max}_{\delta f\in\mathcal{H}_{t}}\Delta L_{f_{t}}[\delta f]\,,$ (6)
where $\mathcal{H}_{t}$ is an appropriate space of functions (satisfying $\mathbb{E}_{\mathbf{x}\sim q_{f_{t}}}\delta f(\mathbf{x})=0$ if Equation 5 is used). The solution to this problem can be interpreted as a Newton step in the space of energy functions. Because, for an EBM, the Fisher information matrix for the energy function and the Hessian of the expected log-likelihood are the same, we can also interpret this solution as a natural gradient step (see Appendix A). This approach is analogous to the second-order step implemented in modern discriminative gradient boosting libraries such as XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017) and which can be traced back to Friedman et al. (2000).
In updating the current iterate, $f_{t+1}=f_{t}+\alpha_{t}\cdot\delta f^{*}_{t}$, we scale $\delta f^{*}_{t}$ by an additional scalar step-size $\alpha_{t}$. This can be interpreted as a globalization strategy to account for the fact that the quadratic approximation in Equation 4 is not necessarily valid over large steps in function space. A common strategy in nonlinear optimization would be to select $\alpha_{t}$ via a line search based on the original log-likelihood. Common practice in discriminative boosting, however, is to interpret this step-size as a regularization parameter and to select a fixed value in ]0,1]]0,1]] 0 , 1 ], with (more) smaller steps typically outperforming fewer larger ones when it comes to generalization. We choose to adopt a hybrid strategy, first selecting an optimal step-size by line search and then shrinking it by a fixed factor. We find that this typically accelerates convergence, allowing the algorithm to take comparatively larger steps that increase the likelihood in the initial phase of boosting.
For a starting point, $f_{0}$, we can choose the logarithm of any probability distribution over $\mathcal{X}$, as long as it is easy to evaluate. Sensible choices are a uniform distribution (i.e., $f\equiv 0$), the product of marginals for the training set or any mixture distribution between these two. In Figure 2 we show an example of NRGBoost starting from a uniform distribution and learning a toy 2D data density.
@@ 3.1 Weak Learners
Figure 2: Density learned by NRGBoost at different boosting iterations (1, 3, 10 and 100), starting from a uniform distribution. The data distribution is depicted on the right (see Appendix D.1 for details). Weak learners are piecewise constant functions given by binary trees with 16 leaves.
As a weak learner we will consider piecewise constant functions defined by binary trees over the input space. Letting $\bigcup_{j=1}^{J}X_{j}=\mathcal{X}$ be the partitioning of the input space induced by the leaves of a binary tree whose internal nodes represent a split along one dimension into two disjoint partitions, we take $\mathcal{H}$ as the set of functions such as
$\delta f(\mathbf{x})=\sum_{j=1}^{J}w_{j}\mathbf{1}_{X_{j}}(\mathbf{x})\,.$ (7)
Here, $\mathbf{1}_{X}$ denotes the indicator function of a subset $X$ and $w_{j}$ are values associated with each leaf $j\in[1..J]$. In a standard decision tree, these values would typically encode an estimate of $p(y|\mathbf{x}\in X_{j})$, with $y$ being a special target variable that is never considered for splitting. In our generative approach, they encode unconditional densities (or more precisely energies) over each leaf’s support and every variable can be used for splitting. Note that the $\delta f$ functions are thus parametrized by the values, $w_{j}$, as well as the structure of the tree and the variables and thresholds for the split at each node, which ultimately determine the $X_{j}$. We omit these dependencies for brevity.
By substituting the definition from Equation 7 into our objective (Equation 5) we get the following optimization problem to find the optimal binary tree:
@@
where $P(X_{j})$ and $Q_{f}(X_{j})$ denote the probability of the event $\mathbf{x}\in X_{j}$ under the respective distribution and the constraint ensures that $\delta f$ has zero expectation under $q_{f}$. With respect to the leaf weights, this is a quadratic program whose optimal solution and objective values are respectively given by
@@
Because carrying out the maximization of this optimal value over the tree structure that determines the $X_{j}$ is hard, we approximate its solution by greedily growing a tree that maximizes it when considering how to split each node individually. A parent leaf with support $X_{P}$ is thus split into two child leaves, with disjoint support $X_{L}\cup X_{R}=X_{P}$, so as to maximize over all possible partitionings along a single dimension, $\mathcal{P}\left(X_{P}\right)$, the following objective:
@@
Note that, when using parametric weak learners, computing a second-order step would typically involve solving a linear system with a full Hessian. As we can see, this is not the case when the weak learners are binary trees, where the optimal value to assign to a leaf $j$ does not depend on any information from other leaves and, likewise, the optimal objective value is a sum of terms, each depending only on information from a single leaf. This would not have been the case had we tried to optimize the log-likelihood functional in Equation 3 directly, instead of its quadratic approximation.
@@ 3.2 Amortized Sampling
To compute the leaf values in Equation 9 and the splitting criterion in Equation 10 we would have to know $P(X)$ and be able to compute $Q_{f}(X)$, which is infeasible due to the intractable normalization constant. We therefore estimate these quantities, with recourse to empirical data for $P(X)$ and to samples approximately drawn from the model with MCMC. Because, even if the input space is not partially discrete, $f$ is still discontinuous and constant almost everywhere, we can not use gradient based samplers. We therefore rely on Gibbs sampling, which only requires evaluating each $f_{t}$ along one dimension at a time, while keeping all others fixed. Importantly, this can be computed efficiently for a tree by traversing it only once. However, since our energy function at boosting iteration $t$ is given by a sum of $t$ trees, this computation scales linearly with the iteration number. As a result, the total computational cost of sampling, over the course of training a model, scales quadratically in the number of boosting iterations, thus precluding us from training models with a large number of trees.
@@
In order to reduce the burden associated with this sampling, which dominates the runtime of training the model, we propose a new sampling approach that leverages the cumulative nature of boosting. The intuition behind this approach is that the set of samples used in the previous boosting round are (approximately) drawn from a distribution that is already close to the current model distribution. It could, therefore, be helpful to keep some of those samples, especially those that conform best to the current model. Rejection sampling allows us to do just that. The boosting update in terms of the densities takes the following multiplicative form:
$q_{t+1}(\mathbf{x})=k_{t}\,q_{t}(\mathbf{x})\exp\left(\alpha_{t}\delta f_{t}(\mathbf{x})\right)\,.$ (11)
@@
Here, $k_{t}$ is an unknown multiplicative constant and since $\delta f_{t}$ is given by a tree, we can easily bound the exponential factor by finding the leaf with the largest value. To sample from $q_{t+1}$, we can therefore use the previous model, $q_{t}$, as a proposal distribution for which we already have a set of samples. We retain each of these previous samples, $\mathbf{x}$, with an acceptance probability of
$p_{accept}(\mathbf{x})=\exp\left[\alpha_{t}\left(\delta f_{t}(\mathbf{x})-\max_{\mathbf{x}}\delta f_{t}(\mathbf{x})\right)\right]\,,$ (12)
noting that knowledge of the normalizing constant, $k_{t}$, is not necessary for its computation.
@@
Our proposed sampling strategy is to maintain a fixed-size pool of approximate samples from the model. At the end of each boosting round, we use rejection sampling to remove samples from the pool and draw new samples from the updated model using Gibbs sampling. Figure 1 depicts a representation of the training loop with this amortized sampling approach. Note that $q_{0}$ is typically a simple model for which we can both directly evaluate the desired quantities (i.e., $Q_{0}(X)$ for a given partition $X$) and efficiently draw exact samples. As such, no samples are required for the first round of boosting. For the second round, we can initialize the sample pool by drawing exact samples from $q_{1}$ with rejection sampling, using $q_{0}$ as a proposal distribution.
@@
This amortized sampling approach works better when either the range of $\delta f_{t}$ or the step sizes $\alpha_{t}$ are small, as this leads to larger acceptance probabilities. Note that, in practice, it can be helpful to independently refresh a fixed fraction of samples, $p_{refresh}$, at each round of boosting in order to encourage a more diverse set of samples between rounds. This can be accomplished by keeping each sample with a probability $p_{accept}(\mathbf{x})(1-p_{refresh})$ instead. Please refer to Algorithm 1 in Appendix A.3 for a high level algorithmic description of the training process with amortized sampling.